### v1

# Atomic Smart Contract Interoperability with High Efficiency via Cross-Chain Integrated Execution

Abstract
With the development of Ethereum, numerous blockchains compatible with Ethereum’s execution environment (i.e., Ethereum Virtual Machine, EVM) have emerged. Developers can leverage smart contracts to run various complex decentralized applications on top of blockchains. However, the increasing number of EVM-compatible blockchains has introduced significant challenges in cross-chain interoperability, particularly in ensuring efficiency and atomicity for the whole cross-chain application. Existing solutions are either limited in guaranteeing overall atomicity for the cross-chain application, or inefficient due to the need for multiple rounds of cross-chain smart contract execution.
To address this gap, we propose IntegrateX, an efficient cross-chain interoperability system that ensures the overall atomicity of cross-chain smart contract invocations. The core idea is to deploy the logic required for cross-chain execution onto a single blockchain, where it can be executed in an integrated manner. This allows cross-chain applications to perform all cross-chain logic efficiently within the same blockchain. IntegrateX consists of a cross-chain smart contract deployment protocol and a cross-chain smart contract integrated execution protocol. The former achieves efficient and secure cross-chain deployment by decoupling smart contract logic from state, and employing an off-chain cross-chain deployment mechanism combined with on-chain cross-chain verification. The latter ensures atomicity of cross-chain invocations through a 2PC-based mechanism, and enhances performance through transaction aggregation and fine-grained state lock. We implement a prototype of IntegrateX. Extensive experiments demonstrate that it reduces up to 61.2% latency compared to the state-of-the-art baseline while maintaining low gas consumption.
Index Terms:
Blockchain Interoperability, Cross-Chain Integrated Execution, Efficient and Atomic Interoperability Protocol, Cross-Chain Smart Contract Invocation.
I Introduction
With the advent of Bitcoin and Ethereum [1, 2], we have witnessed the emergence of an increasing number of programmable blockchains  [3, 4, 5, 6]. On these programmable blockchains, developers can write and deploy smart contracts to build various complex decentralized applications (dApps, such as DeFi, NFT, etc. [7, 8]). Among these programmable blockchains, those that share compatible smart contract execution environment with Ethereum (i.e., Ethereum Virtual Machine, a.k.a. EVM [9]) dominate the landscape, accounting for over 90% of the total value locked [10]. With the increasing number of EVM-compatible blockchains and the diversity of dApps running on each chain, the demand for cross-chain dApps has grown significantly [11]. Cross-chain dApps refer to dApps that require cross-chain smart contract invocations (CCSCI) and coordinated execution across multiple blockchains [12].
However, ensuring overall atomicity for the entire cross-chain dApp while efficiently handling CCSCI remains a critical challenge. As illustrated in Figure 1 left, consider a classic train-and-hotel problem [13]. A user wants to book an outbound train ticket (on Train Chain) through a travel agency (on Agency Chain), then book a hotel (on Hotel Chain), followed by a return train ticket (again on Train Chain). The user wants to ensure that the entire series of CCSCI either all succeed or all fail, ensuring overall atomicity. More importantly, maintaining efficiency during CCSCI is crucial. Key considerations include how to minimize latency, reduce gas consumption (monetary cost), and improve concurrency during the CCSCI process. More cross-chain dApp scenarios are discussed in Section VIII.
Existing CCSCI interoperability solutions generally either fail to ensure the overall atomicity of a cross-chain dApp or guarantee overall atomicity but with low efficiency. To ensure atomicity in the CCSCI process, a widely adopted approach is to use a two-phase commit (2PC) mechanism [14, 12] involving locking and unlocking. For example, some works [15, 16, 17, 18, 19, 20, 21, 22] propose general cross-chain communication protocols that facilitate the transfer of information and data between blockchains through bridging smart contracts deployed on each blockchain. However, these approaches typically ensure at best the atomicity of single-step cross-chain interactions (Figure 1, a single arrow), but fail to guarantee overall atomicity for the entire cross-chain dApp. Some other recent studies attempt to explore how to ensure overall atomicity for the cross-chain dApp [23, 24, 25, 12]. To ensure overall atomicity, the relevant states must remain locked throughout the entire CCSCI process. However, they commonly face challenges in achieving efficiency. The main reason is that, these approaches usually require multiple rounds of cross-chain execution and cross-chain information transfer when handling a cross-chain dApp (Figure 1, left), since a cross-chain dApp usually involves interdependent execution logic distributed over multiple blockchains. It is evident that this method tends to be time-consuming and inefficient (as the locking time could be long), especially when dealing with complex CCSCI. More related work (e.g., cross-chain asset swap, smart contract portability) is discussed and differentiated in detail in Section II.
To fill the research gap, we propose IntegrateX, an efficient interoperability system that guarantees overall atomicity for the cross-chain dApp across EVM-compatible blockchains. To enhance the efficiency of CCSCI, our core idea is to clone and deploy [^1] The logic on the original chain still exists and continues to function normally. the logic of all contracts involved in a cross-chain dApp—originally distributed across multiple chains—onto a single blockchain. This chain thus integrates the entire execution logic of the cross-chain dApp. When the CCSCI is required, this chain can perform integrated execution in one transaction for all related logic, after receiving the necessary states (Figure 1, right). Since multi-round cross-chain executions and interactions are no longer required, IntegrateX maintains high efficiency, even in complex cross-chain dApps.
Figure 1: An example of existing CCSCI solutions (left) and IntegrateX (right) in the Train-and-Hotel scenario.
The design of IntegrateX faces several challenges, which we will address below.
Challenge 1: How to efficiently and securely deploy smart contracts across chains to the same blockchain. We address this by proposing an on-chain/off-chain Hybrid Cross-Chain Smart Contract Deployment Protocol. Specifically, to reduce the overhead associated with cross-chain deployment of smart contracts (thereby improving efficiency), we devise a set of guidelines that allow developers to decouple smart contracts into logic execution contracts and state storage contracts. This approach enables our protocol to clone and deploy only the logic contracts onto the same target blockchain, while the state-heavy contracts remain in their original locations. To further reduce the gas consumption during cross-chain deployment, we propose to clone and deploy the logic contracts to the same blockchain via an off-chain mechanism. However, malicious actors may tamper with the contracts during the off-chain clone and deployment process, leading to compromised security. To address this, we design and deploy bridging smart contracts on each blockchain. These bridging smart contracts perform on-chain cross-chain comparison and verification of the logic contracts between the source chain and the target chain to ensure security.
Challenge 2: How to enhance concurrency and reduce overhead during CCSCI while ensuring overall atomicity. We address this by proposing a Cross-Chain Smart Contract Integrated Execution Protocol. Specifically, to ensure overall atomicity during CCSCI, we adopt a 2PC-based mechanism, similar to existing mainstream approaches. This process involves locking the relevant states across the involved chains, transmitting the states to the chain responsible for integrated execution, executing all logic, and then returning the states to the respective chains to unlock and update them. This guarantees that the cross-chain dApp either completes entirely or is fully rolled back. Moreover, cross-chain state transfers may incur significant overhead. To mitigate the overhead of cross-chain communication, we aggregate multiple cross-chain transactions into a single one where necessary, and transmit them across chains. More importantly, a straightforward state-locking mechanism reduces transaction concurrency. To enhance concurrency, we establish a set of guidelines that allow developers to further decompose and lock contract states at a finer granularity. By locking more granular states, our protocol alleviates the issue of poor concurrency that arises when entire states are locked.
This paper mainly makes the following contributions:
  • We present IntegrateX, a cross-chain interoperability system that efficiently facilitates CCSCI while ensuring overall atomicity for the cross-chain dApp. IntegrateX can be flexibly deployed on EVM-compatible blockchains without requiring modifications to the underlying blockchain systems.
  • In IntegrateX, we propose the Hybrid Cross-Chain Smart Contract Deployment Protocol. It achieves efficient and secure cross-chain deployment through the decoupling of smart contract logic and state, and the hybrid approach of off-chain logic deployment and on-chain comparison verification.
  • In IntegrateX, we propose the Cross-Chain Smart Contract Integrated Execution Protocol. It ensures overall atomicity of cross-chain invocations through a 2PC-based atomic integrated execution mechanism, and enhances the protocol efficiency through an aggregated cross-chain transaction transmission mechanism and fine-grained state lock.
  • We implement a prototype of IntegrateX and make it open source [26]. Extensive experiments based on real-world use cases demonstrate that IntegrateX reduces latency by up to 61.2% meanwhile maintains low gas cost and high concurrency compared to the state-of-the-art baseline. In more complex cross-chain invocations, IntegrateX will further improve efficiency.
II Background and Related Work
Cross-Chain Asset Swap and Transfer. Blockchain interoperability has gained significant attention. One widely studied area is cross-chain asset swap and transfer [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. Cross-chain asset swap protocols enable untrusted parties to exchange assets across different blockchain networks, while cross-chain asset transfer protocols involve locking or burning assets on the source chain and creating their equivalents on the target chain. While these protocols provide atomicity guarantees during cross-chain asset swaps or transfers, their protocols cannot be applied to our target scenario of CCSCI.
Interoperability Protocol for General Communication. Besides cross-chain asset transfers and swaps, some other efforts aim to provide interoperability protocols for more general cross-chain communication [15, 16, 17, 18, 19, 20, 21, 22], enabling the transfer of general data beyond assets between blockchains. However, the primary limitation of these protocols is that they typically guarantee atomicity for, at most, a single cross-chain step between two blockchains. For general cross-chain dApps, which often require multiple rounds of cross-chain interactions across multiple chains, these protocols fail to ensure atomicity for the entire cross-chain dApp. In contrast to these approaches, IntegrateX guarantees overall atomicity of the entire multi-round CCSCI for general cross-chain dApps.
CCSCI with Overall Atomicity. There are some studies that attempt to ensure overall atomicity for the cross-chain dApp with complex CCSCI [38, 39, 40, 41, 42, 23, 24, 25]. However, these approaches have various limitations. For example, HyperService [38] only can provide weak atomicity, named financial atomicity, via an insurance mechanism. It fails to provide strong execution atomicity guarantees during CCSCI. Some works like Heterogeneous Paxos [42] provide overall atomicity for CCSCI. However, they usually require the underlying blockchains to make targeted modifications to fit their protocols, limiting their generality. Some other works, like Unity [39], also ensure atomicity for a whole CCSCI process. However, their protocols only apply to cross-chain scenarios between two blockchains and does not take into account the more general multi-chain case.
There are few other solutions that guarantee overall atomicity for the cross-chain dApp with complex CCSCI. However, they generally suffer from inefficiency  [23, 24, 25]. A complex cross-chain dApp typically involves multiple blockchains and several interdependent application logic components (implemented through smart contracts). In these works, handling such complexity requires multiple rounds of cross-chain execution and cross-chain interaction in sequential order across several involved chains. Specifically, this involves locking relevant states, executing the logic fragment, reaching consensus, transferring intermediate states to the next blockchain responsible for executing the subsequent logic fragment, and finally unlocking and updating the states. As a result, they typically suffer from high latency and poor concurrency handling CCSCI. Additionally, Atomic IBC [24] relies on the Cosmos Hub to provide security guarantees for its protocol. This approach is not easily adaptable to other ecosystems (e.g., EVM-compatible blockchains) and may introduce security bottlenecks.
Unlike these aforementioned works, IntegrateX is able to maintain high efficiency during CCSCI while ensuring overall atomicity. Moreover, it can be deployed to blockchains with the same execution environment (e.g., EVM-compatible) without requiring modifications to the underlying blockchain.
Smart Contract Portability. Another type of approach, known as smart contract portability [43, 44, 45], involves frequently moving or replicating entire smart contracts with all the states associated across different blockchains. This allows cross-chain invocations to be converted into intra-chain invocations during contract execution. However, the practicality of this approach is often questionable. The choice of which blockchain to deploy and run a smart contract and dApp typically involves several considerations, including security, ecosystem, and business concern. Moving an entire smart contract from one ecosystem to another is often not favored by developers, as it may compromise security or disrupt the completeness of the ecosystem. Besides, smart contracts often manage a significant amount of user state, and frequently relocating all of this state across chains incurs substantial overhead. This inefficiency further reduces the practicality of these approaches.
The philosophy of IntegrateX is fundamentally different from smart contract portability. In IntegrateX, the states involved in each smart contract are still maintained and updated by their original blockchains. This allows dApps to continue benefiting from the security and ecosystem of their original blockchain for most of the time. During the CCSCI process, we only temporarily lock part of the state relevant to the invocation and transfer it to a single blockchain, where the logic is integrated and executed using that blockchain’s execution environment.
III System Overview
In this section, we provide an overview of IntegrateX. IntegrateX’s architecture and workflow overview is shown in Figure 2. For a detailed analysis of IntegrateX’s security and further discussion, please refer to Section VI and Section VIII.
III-A System and Threat Model
In IntegrateX, there are $n$ (variable) blockchains that share the same smart contract execution environment (e.g., EVM). These blockchains are developed by their respective projects (e.g., Ethereum, BSC, etc. [2, 46]) and operated by their own blockchain nodes, which handle transaction processing, consensus within the blockchain, and other tasks. IntegrateX also features $m$ (variable) relayers responsible for trustless cross-chain communication between blockchains, similar to many mainstream interoperability protocols [24, 17, 47, 48]. Each relayer has its own public-private key pair and signs cross-chain transactions. Additionally, IntegrateX’s bridging smart contracts are deployed on each blockchain (similar to on-chain light clients [17, 24]). The bridging contracts mainly serve to verify state transitions on the other blockchains, and send information externally via event emissions. The above IntegrateX’s transport layer architecture is similar to some well-known cross-chain communication protocols (e.g., IBC) [17, 24], which guarantee basic security during cross-chain communication. However, IntegrateX achieves efficiency and overall atomicity at the application layer, which these other protocols do not.
IntegrateX also includes several intra-chain dApp providers. These providers can freely choose to deploy their contracts and run their intra-chain dApps on different blockchains. Additionally, there are cross-chain dApp providers who can similarly flexibly choose to deploy their cross-chain dApps across various blockchains. A cross-chain dApp typically consists of multiple intra-chain dApps distributed across different blockchains, and relies on IntegrateX’s cross-chain protocol to ensure efficient CCSCI and maintain the overall atomicity of the cross-chain dApp. Finally, there are also users within the IntegrateX ecosystem. These users interact with the cross-chain dApps by sending transactions to the blockchains in order to use these cross-chain dApps.
Threat Model. In IntegrateX, we make only minimal trust assumptions about the relayers, assuming that at least one relayer is honest and functioning correctly. This assumption is consistent with those made in many existing secure interoperability protocols [24, 17, 47, 48]. For each blockchain, the proportion of Byzantine nodes is assumed to be less than the fault tolerance threshold $t$ of the respective blockchain network (e.g., for blockchains using BFT-type consensus protocols under partial-synchronous network, $t=1/3$ [49]). This ensures the safety and liveness of consensus within each blockchain. As for the dApp providers and users—who represent the application layer components—we make no specific threat assumptions, similar to most existing works [17, 25, 23]. However, we do discuss common countermeasures for dealing with malicious behavior from these components in Section VIII.
III-B Objective
IntegrateX aims to achieve the following primary objectives:
  • Efficiency: During the process of cross-chain smart contract deployment and invocation, IntegrateX seeks to reduce latency, lower gas consumption, and increase transaction concurrency.
  • Overall Atomicity: IntegrateX aims to ensure overall atomicity for cross-chain dApps that require it during CCSCI. This means guaranteeing that the series of CCSCI operations required by cross-chain dApp providers either all succeed or all fail.
In addition, IntegrateX aims to possess the following desirable properties. Reliability: IntegrateX ensures that cross-chain transactions can still be completed even in the presence of malicious relayers. Verifiability: IntegrateX guarantees that cross-chain transactions can be verified for authenticity, completeness, and validity, even if malicious relayers are involved. Consistency: IntegrateX ensures that the state changes across the blockchains involved in the cross-chain transaction remain coordinated and consistent.
The proofs and experiments related to aforementioned properties are detailed in Section VI and Section VII.
Figure 2: An overview of IntegrateX.
III-C Primary Workflow
The operation of IntegrateX consists of three main phases: Smart Contract Preparation, Cross-Chain Smart Contract (CCSC) Deployment, and Cross-Chain Smart Contract (CCSC) Integrated Execution, as illustrated in Figure 2.
Smart Contract Preparation. This phase only occurs when dApp providers need to develop or upgrade smart contracts, so the frequency of this phase is relatively low. In this phase, dApp providers can flexibly develop logic contracts and state contracts according to our defined logic-state decoupling guidelines (Section IV-A). This decoupling is primarily intended to reduce gas consumption and minimize storage costs on the target chain by cloning only the logic contracts during the subsequent CCSC deployment phase. Additionally, dApp providers can develop their contracts according to our fine-grained state lock guidelines (Section V-C), which enhances transaction concurrency during the CCSC integrated execution phase. Once the smart contract development is complete, dApp providers can freely choose a blockchain to deploy their contracts.
CCSC Deployment. This phase only runs when cross-chain dApp providers need to deploy logic contracts across chains onto the same blockchain, which also occurs infrequently. In this phase, cross-chain dApp providers can flexibly choose a target chain for deployment. By sending a transaction, they can clone and deploy the specified logic contracts from other designated chains onto the chosen target chain based on their cross-chain dApp needs. IntegrateX introduces an off-chain clone approach (Section IV-B) to reduce gas consumption, while on-chain cross-chain verification (Section IV-C) ensures that the cloned contracts are identical to the original ones.
CCSC Integrated Execution. This phase runs every time a user needs to interact with a cross-chain dApp, making it the most frequent phase of operation. In this phase, users interact with cross-chain dApps by sending transactions to the target chain, invoking cross-chain smart contracts based on the dApp’s application logic. IntegrateX employs an atomic integrated execution scheme (Section V-A), similar to the 2PC protocol, to ensure the overall atomicity of the series of CCSCI involved in a cross-chain dApp. This scheme first locks the relevant states of the smart contracts involved in the cross-chain dApp on the respective chains at a fine-grained level, then transmits these states across chains to the target chain. Since the target chain already contains all the execution logic required for the cross-chain dApp (from the earlier phase), it can integrate and execute all logic within one transaction once the necessary states have been received. After execution, the new states are returned to the corresponding chains for unlocking and updating. Additionally, IntegrateX employs a transaction aggregation mechanism (Section V-B) during this phase to reduce cross-chain overhead.
In the subsequent descriptions within this paper, we will refer to the target chain selected by a cross-chain dApp for deployment—i.e., the chain responsible for integrated execution—as the execution chain. The other chains associated with the cross-chain dApp will be collectively referred to as invoked chains.
IV Hybrid Cross-Chain Smart Contract Deployment Protocol
IV-A Logic-State Decoupling
To efficiently perform CCSCI, we need to deploy the logic of the invoked contracts to the same execution chain for efficient integrated execution. Existing contracts often contain both logic and state, and directly cloning such contracts would result in high gas costs and additional state storage overhead on the execution chain. Therefore, we design a set of Logic-State Decoupling (LSD) Guidelines to guide the developers to decouple existing contracts into logic execution contracts and state storage contracts. During cross-chain deployment, only the logic contracts need to be cloned, which reduces gas costs. Developers can follow these guidelines to develop new contracts with separated logic and state or upgrade existing contracts by decoupling them. We now provide a detailed explanation of the LSD Guidelines. Moreover, a simple example illustrating the LSD is provided in Section IV-A1.
State Contract. According to the LSD Guidelines, the decoupled state contract first includes all the variables (representing states) from the original contract, as well as all the view functions that read the contract’s state. Since view functions are read-only and do not generate transactions, they do not affect cross-chain execution. Additionally, the state contract contains functions for locking and updating the contract state, as cross-chain invocations require locking and updating states. The state contract should also contain functions that call the logic contract in order to interact normally with the logic contract.
Logic Contract. In the logic contract, no variables are stored. It only contains the functions that implement the original contract’s execution logic. These functions are called by the state contract to carry out the dApp’s logic operations. When the state contract calls these functions, it passes all necessary state data (variables), and after the functions complete execution, the results are returned to the state contract. In our protocol, only the logic contract needs to be cloned, which reduces gas costs during cross-chain deployment.
Remarks. Our protocol also supports existing smart contracts, even if they are not decoupled into logic and state. However, in such cases, the cross-chain deployment process will incur higher gas costs. In this case, the contracts deployed to the execution chain do not actively maintain their own state. Instead, they passively update their state during each cross-chain invocation. The discussion related to developers’ learning costs is given in Section VIII.
IV-A1 Logic-State Decoupling Example
We now give a sample of logic-state decoupling. In the following code 1, after applying logic-state decoupling, the original Hotel contract is split into two separate contracts: the logic contract LHotel and the state contract SHotel. The LHotel contract contains no state variables and only includes the book() function, which implements the hotel reservation functionality. Since there are no variables within the LHotel contract, the book() function must take all necessary parameters as inputs.
On the other hand, the SHotel contract retains all the variables from the original Hotel contract and introduces an additional address variable, addr_lhotel, which records the address of the LHotel contract. In the book() function of the SHotel contract, no reservation logic is implemented directly; instead, it calls the book() function from LHotel using the addr_lhotel parameter to execute the hotel reservation functionality.
By decoupling the logic and state in this way, only the LHotel contract needs to be cloned. Since LHotel contains no state variables, this approach significantly reduces gas consumption during the cross-chain clone and deployment process.
contract Hotel{
int256 price;
int256 remain;
mapping (address => int256) accounts;
function getPrice() public view returns(uint256);
function getRemain() public view returns(uint256);
function book(address user_addr, uint256 num) public returns(uint256);
function LockState(bytes[] memory args) public returns();
function UpdateSteate(bytes[] memory args) public returns();
contract LHotel{
function book(uint256 price, uint256 remain, uint256 num) public returns(uint256)
contract SHotel{
int256 price;
int256 remain;
address addr_lhotel;
mapping (address => int256) accounts;
function getPrice() public view returns(uint256);
function getRemain() public view returns(uint256);
function LockState(bytes[] memory args) public returns();
function UpdateSteate(bytes[] memory args) public returns();
function book(address user_addr, uint256 num) public returns(uint256);
Listing 1: Pseudocode of Hotel Logic-State Decoupling
(a) Off-chain clone and deployment
(b) On-chain verification
Figure 3: Hybrid Cross-Chain Smart Contract Deployment Protocol.
IV-B Off-Chain Clone and Deployment
Transmitting contract bytecode on-chain (via contract event) requires a significant amount of gas. To improve the efficiency of cross-chain logic deployment and reduce gas consumption, we propose transferring the contract bytecode via an off-chain solution. The off-chain clone and deployment consists of two phases: the preparation phase and the clone phase. The main process is shown in Figure  3(a).
Preparation. In this phase, the cross-chain dApp provider needs to obtain the call tree of smart contracts and determine which logic contracts need to be cloned. This can be achieved by using some static analysis tools [50] and other methods, similar to existing works [51, 23, 25]. After this, the developer can choose a blockchain to deploy the cross-chain dApp. It is important to note that we offer developers a high degree of flexibility: They can select any blockchain according to their preference. To reduce costs, they may also choose a blockchain that already contains some of the invoked logic contracts. Once the selection is made, the developer sends a transaction to invoke the bridging contract on the chosen chain. The bridging contract then triggers an event to notify the relayers to initiate the cross-chain deployment. The event includes the ID of the invoked chain and the addresses of the logic contracts $Addr_{\text{L}}$ to be cloned on the invoked chain. This concludes the preparation phase.
Clone. After the event is triggered, relayers will detect the event and use the getcode() function from the bridging contract on the invoked chain to obtain the bytecode of the contract that needs to be cloned. This process is an off-chain read-only inquiry and, therefore, does not consume gas. Subsequently, the relayers will obtain the ABI file which defines the contract interface. The relayers will then deploy the contract to the execution chain. Once a relayer completes the redeployment, it registers the address of the cloned logic contract $Addr^{\prime}_{\text{L}}$, through the bridging contract on the execution chain, marking the end of the clone phase.
Remarks. For the cross-chain reliability of the protocol, multiple relayers perform the clone process after detecting the bridging contract’s event. Therefore, even if some relayers do not respond to the event, other relayers will ultimately complete the contract clone and deployment. Once one relayer completes the deployment, the bridging contract will trigger an event to stop the others, ensuring that the logic contract will not be deployed multiple times.
IV-C On-Chain Verification
Although off-chain clone and deployment can achieve efficient cross-chain logic deployment, it does not guarantee security as there may be malicious relayers present in the system. A malicious relayer could potentially modify the contract bytecode or deploy a wrong contract. To ensure verifiability and security during the off-chain clone and deployment process, we propose the on-chain cross-chain verification scheme to compare the cloned logic contract with the original contract and verify its correctness. The on-chain verification process is shown in Figure  3(b).
After the cross-chain deployment is completed, the cross-chain dApp provider can initiate cross-chain verification on the execution chain by calling the Verification() function of the bridging contract. The bridging contract will search the cloned contract bytecode of address $Addr^{\prime}_{\text{L}}$, and calculate the hash of the bytecode. Then the bridging contract triggers an event includes the hash computation result. The relayers are responsible for transmitting this information (a cross-chain transaction) along with its Merkle proof [52] to the invoked chain. In the invoked chain, the Merkle proof with the cross-chain transaction is first verified to ensure the result has already reached consensus on the execution chain. The bridging contract then searches the corresponding local contract bytecode based on the address $Addr_{\text{L}}$, and calculates the hash of the bytecode. Then, the bridging contract will verify whether it matches the hash transmitted across the chain. The result of the verification is then returned.
If the verification is successful, the cloned contract will be marked as verified and allowed for subsequent cross-chain invocations, and the relayer that performed the cross-chain deployment will be rewarded. If the verification fails, the relayer will be penalized, and the off-chain clone and deployment process will be restart.
Remarks. For the reliability of the cross-chain protocol, multiple relayers could transmit the same cross-chain transaction. However, the bridging contracts will deduplicate identical transactions from multiple relayers to avoid multiple executions on-chain. This process also applies in the subsequent integrated execution protocol.
V Cross-Chain Smart Contract Integrated Execution Protocol
V-A Atomic Integrated Execution
To efficiently and atomically execute complex CCSCI, we propose an atomic integrated execution scheme. As all the invoked logic has been migrated onto the execution chain, the atomic integrated execution does not need multiple rounds of cross-chain execution when handling CCSCI. This enables all the logic to be executed within a single transaction, enhancing the efficiency of CCSCI. Moreover, to ensure overall atomicity for a series of CCSCI, we employ a state synchronization mechanism based on the Two-Phase Commit protocol [14]. This process locks the relevant states across the involved chains, transmits the states to the chain responsible for integrated execution, and then returns the states to the respective chains to unlock and update them after the execution is completed. The entire atomic integrated execution consists Locking, Integrated Execution and Updating, which is shown in the Figure 4.
Locking. The Locking process locks all the required invoked contract states on the invoked chain. The cross-chain dApp provider can obtain all the required state on the call tree beforehand via tools such as static analysis [50]. Based on this information, a user sends a transaction via the cross-chain dApp to invoke the cross-chain dApp contract. The cross-chain dApp contract then calls the bridging contract to issue an event to lock the relevant states on the invoked chains. When relayers detect the event, it will transfer this message (i.e., cross-chain transaction with Merkle proof) to each invoked chain. The bridging contract on the invoked chain will verify the authenticity of the cross-chain though calculating the Merkle proof of the transaction and invoke the LockState() function of each invoked contract. Once the bridging contract has successfully locked the state and retrieved the required contract states, it will trigger an event to return the states. After the relayers detect the event, they will transmit these states via cross-chain transactions to the execution chain. After the execution chain’s bridging contract verifies the authenticity of the transactions via the Merkle proof, it will return the states to the dApp. Once all requested states are returned from individual invoked chains, the Locking process ends, and these states are used as inputs for the Integrated Execution.
Integrated Execution. The Integrate Execution process executes the entire CCSCI logic on the execution chain. Cross-chain dApp contracts on the execution chain use the requested state values as inputs to perform the full call tree execution. Since all contracts required for the cross-chain invocation have completed logic migration and have been verified already, the integrated execution can be completed within a single transaction on the execution chain. The cross-chain dApp contract records the output results of each invoked contract during the Integrated Execution, allowing for state updates of the invoked contracts on other chains after the execution is completed. Once execution is complete, the Integrated Execution ends and transitions to the Updating process.
Updating. The Updating process unlocks and updates all the invoked contract states on the invoked chains. After Integrated Execution is completed on the execution chain, the cross-chain dApp contract triggers an event to update the result via bridging contract. The relayers will distribute the result to the invoked chains. The bridging contract on each invoked chain verifies the Merkle proof of the cross-chain transaction and then invokes the UpdateState() function of each invoked contract, which will unlock and update the states of the invoked contracts.
Remarks: Rollback. During the Atomic Integrated Execution, state rollback may occur due to failure in locking the state or execution failure. The invoked state might already be locked by another cross-chain invocation, which causes the failure to lock the invoked state. In this case, the bridging contract on the execution chain will initiate a new event to unlock all associated contracts that have already been locked and returned in this cross-chain invocation. For the contracts that have not yet completed the locking process, the event will cancel the lock attempt, thereby ensuring overall atomicity. The execution might also fail, due to the reason such as insufficient gas fee or insufficient states. In this case, the bridging contract on the execution chain will discard the obtained states and initiate a cross-chain event to unlock all locked contracts, thus ensuring all invoked contracts are either fully locked or not locked at all.
Figure 4: The CCSCI process in sequential invocation (left) and in IntegrateX’s Cross-Chain Smart Contract Integrated Execution Protocol (right).
Timeout. Additionally, to prevent the invoked contract state from being locked for an extended period, a design of timeout will be determined by the dApp developer within the application and managed by the execution chain. For transactions that time out, such as when the locking is not completed or execution is not finished for an extended period, the execution chain will mark the cross-chain call as failed and send a cross-chain event to unlock the relevant contracts.
Finality. In IntegrateX, all cross-chain transactions must wait until the consensus on the initiating chain is finalized (or, for Nakamoto-type consensus, highly likely to be finalized) before being committed to the target chain, to ensure cross-chain security.
V-B Transaction Aggregation
A large amount of cross-chain transactions for transmitting state incurs significant gas consumption. As shown in Figure 4 left, handling CCSCI by sequential invocation may need to invoke contracts on the same chain in multiple rounds, which cause multiple rounds of cross-chain state transfers. To address this issue, we design a transaction aggregation mechanism to reduce gas costs caused by multiple invocations of different contracts and state transfers on the same chain.
As shown in Figure 4 right, our protocol locks all required states simultaneously, allowing multiple states on the same chain to be locked together, even when the calls are non-contiguous. The transaction aggregation mechanism combines all state requests on the same chain into a single transaction, reducing the number of cross-chain transactions. Similarly, during state updates, the transaction aggregation mechanism reduces the number of update requests. Therefore, for cross-chain calls involving multiple contracts on the same chain, this approach ensures that the number of cross-chain transactions is equal to the number of invoked chains (rather than the number of contracts), thereby reducing gas consumption.
V-C Fine-Grained State Lock
During the atomic integrate execution process, we need to lock contract states to ensure atomicity. A simple approach is to either lock the entire state of the invoked contract or lock individual states being invoked [25, 23]. However, these state-locking mechanism reduces transaction concurrency. Because once a state is locked, any subsequent transactions related to that state will fail until the state is unlocked. Therefore, we establish a set of guidelines to guide developers in decomposing certain states that can be split into finer granularity, and allowing partial state locking. Unlike existing protocols that require locking the entire state, our fine-grained state lock mechanism locks only partial of a state at a fine-grained level. This approach enhances concurrency by reducing unnecessary state locking.
In EVM-compatible blockchains, Solidity is the smart contract language. In Solidity, the state of a contract is typically represented by variables. Various types of variables are used in Solidity, such as uint, address, and boolean. We find that the uint variable is widely used and can be decomposed. Based on this observation, we design a fine-grained state lock specifically for uint variables and develop a lock pool mechanism. The lock pool is a structure where, during the use of the fine-grained state lock, part of the state is locked within this structure until execution is completed, while the unlocked portion of this state remains accessible, thereby enhancing the concurrency of the application.
For variables that can be directly derived from transaction inputs, the exact value of the required state can be precisely determined. A fine-grained state lock can be applied to accurately lock only the relevant portion of such state. For states that are dynamically used during execution, their exact values cannot be determined at the beginning. We allow dApp developers to lock these states in fixed-size increments based on their needs.
Remarks. We focus on the flexibility of this mechanism, allowing developers to choose whether to implement the finer granularity locking mechanism in their dApp and to set the lock granularity based on the specific needs of the dApp. Moreover, developers can set the fixed size of the fine-grained state lock based on their preferences, tailored to the specific use case of the application. More discussion related to developers’ learning costs is given in Section VIII.
VI Security Analysis
VI-A Security in Hybrid Cross-Chain Smart Contract Deployment Protocol
Theorem 1.
The Hybrid Cross-Chain Smart Contract Deployment Protocol ensures reliability, verifiability, and consistency, if the proportion of Byzantine nodes in each blockchain is less than its fault tolerance threshold, and at least one functional relayer is present.
Proof.
Reliability. The Hybrid Cross-Chain Smart Contract Deployment Protocol ensures reliability, meaning that when an execution chain issues a request for cross-chain deployment of a contract, the requested contract will eventually be deployed to the execution chain. Within the Hybrid Cross-Chain Smart Contract Deployment Protocol, multiple relayers listen for such requests. Even if malicious relayers intentionally ignore the requests, assuming that at least one functional relayer exists, this relayer will ultimately handle the logic clone and deployment of the requested contract. Thus, even in the worst-case scenario, the protocol ensures that the contract will be successfully deployed to the execution chain.
Verifiability. The Hybrid Cross-Chain Smart Contract Deployment Protocol ensures verifiability, meaning that both the execution chain and the invoked chain are able to verify the cross-chain transactions transmitted by the relayers. Additionally, both chains can verify the hash values of the contract bytecode before and after the cross-chain clone and deployment to ensure that the contract has been deployed correctly.
In the Hybrid Cross-Chain Smart Contract Deployment Protocol, multiple relayers listen to the request and relay the messages. The invoked chain (bridging contract) validates the authenticity of the transactions using the Merkle proof attached with the cross-chain transactions, thereby preventing the relayers from altering the cross-chain data. By parsing the transactions, the invoked chain can obtain the bytecode hash of the cloned contract and compare it with the bytecode hash of the original contract on the chain to ensure the correctness of the cross-chain deployment. Additionally, the nonce value associated with each transaction prevents malicious replay attacks.
Consistency. The Hybrid Cross-Chain Smart Contract Deployment Protocol ensures consistency, meaning that during the off-chain clone and deployment as well as the on-chain verification process, both the execution chain and the invoked chain reach agreement on the outcome of the cross-chain requests. In the Hybrid Cross-Chain Smart Contract Deployment Protocol, the proportion of malicious nodes on both the execution chain and the invoked chain remains within the fault tolerance threshold. Moreover, the protocol requires to wait until consensus on one chain is finalized (or highly likely to be finalized) before committing the cross-chain transaction to another chain. As a result, even in the presence of Byzantine nodes, each chain can still achieve consensus on cross-chain transactions and finalize them, ensuring that consistency is not undermined by malicious nodes. This guarantees that the outcomes of cross-chain requests remain consistent. ∎
VI-B Security in Cross-Chain Smart Contract Integrated Execution Protocol
Theorem 2.
The Hybrid Cross-Chain Smart Contract Integrated Execution Protocol ensures overall atomicity, reliability, verifiability, and consistency, if the proportion of Byzantine nodes in each blockchain is less than its fault tolerance threshold, and at least one functional relayer is present.
Proof.
Overall Atomicity. The Cross-Chain Smart Contract Integrated Execution Protocol guarantees overall atomicity, which means that in the selected CCSCI process, state changes on both the execution chain and the invoked chain either all succeed or all fail, preventing any situation where one chain’s state changes while the other does not. We employ the atomic integrated execution mechanism, similar to the 2PC scheme. For one cross-chain dApp, during this process, all the states required by the invocation on the invoked chains will be locked. If any contract has already been locked by another invocation, this invocation will fail, and all other invoked contracts will be unlocked. If the execution chain obtains all the necessary states but the execution fails due to insufficient gas or other reasons, the execution will be aborted, and all related locked contracts will be unlocked without any state changes. Furthermore, the protocol incorporates a timeout scheme: If any of the invoked chains fails to return the required state within the specified time frame by dApp, or if the execution transaction on the execution chain fails to complete within the specified time limit, the execution chain will abort the invocation and unlock all related contract states, ignoring any subsequent state responses from the invoked chains. As a result, only when the execution chain has successfully acquired all required states and completed execution will it issue state updates to all related invoked chains, thereby ensuring the atomicity of the entire CCSCI process.
Reliability. The Cross-Chain Smart Contract Integrated Execution Protocol ensures reliability, which means that when an execution chain initiates a CCSCI request, the invoked chain will eventually receive the cross-chain transaction, and the state returned by the invoked chain will likewise be received by the execution chain. In the Cross-Chain Smart Contract Integrated Execution Protocol, multiple relayers monitor CCSCI requests. Even in the presence of malicious relayers who deliberately fail to respond to the request, the assumption of at least one functional relayer ensures that, in the worst-case scenario, this relayer will relay the cross-chain transaction to the invoked chain, ensuring that the transaction is eventually received. Similarly, even in the worst-case scenario, at least one relayer will transmit the state returned by the invoked chain back to the execution chain, thereby guaranteeing the reliability of the CCSCI process.
Verifiability. The Cross-Chain Smart Contract Integrated Execution Protocol ensures verifiability, which means the ability of the invoked chain to verify the authenticity of cross-chain transactions transmitted by relayers from the execution chain, while the execution chain can also verify the transactions returned by the invoked chain. In this protocol, multiple relayers listen for the request and relay messages. Both the execution chain and the invoked chain can independently validate the authenticity of the cross-chain transactions using the Merkle proof attached with the transactions. Additionally, the use of transaction nonce values prevents malicious replay attacks, ensuring the integrity of the cross-chain interaction.
Consistency. The Cross-Chain Smart Contract Integrated Execution Protocol guarantees consistency, ensuring that both the execution chain and the invoked chain agree on the result of the requested operation during a cross-chain smart contract invocation. In this protocol, the proportion of malicious nodes on both chains remains within the fault tolerance threshold, allowing consensus to be achieved even if there exist Byzantine nodes. Additionally, the protocol requires waiting until consensus on one chain has been finalized before committing the cross-chain transaction to the other chain. Therefore, cross-chain transactions on the blockchain cannot be maliciously altered, ensuring that both the execution chain and the invoked chain reach a unified agreement on the outcome of cross-chain operations. ∎
VII Implementation and Evaluation
VII-A Implementation
We implement a prototype of IntegrateX based on a cross-chain communication project Bitxhub [53] (its transport layer is similar to IBC). The relayers are implemented in Golang, while the bridging contracts are implemented using Solidity. In addition, we implement several EVM-compatible blockchains in Golang leveraging go-ethereum clients. For performance comparison, we also implement a baseline CCSCI protocol GPACT [23]. As mentioned in the paper, GPACT is one of the most advanced existing protocols that can guarantee overall atomicity for complex CCSCI. However, GPACT has limitations in efficiency, as it requires multiple rounds of cross-chain execution and message exchange. Furthermore, we use Solidity to implement the Train-and-Hotel example mentioned in the paper, along with smart contracts in various other scenarios, to demonstrate the IntegrateX’s performance in real use cases and under different conditions.
We test the performance of IntegrateX on multiple servers. Each server is equipped with an Intel® Core(TM) i5-10400F CPU @ 2.90 GHz and 15.5 GB RAM running 64-bit Ubuntu 22.04 LTS. We deploy 4 relayers via Docker containers on multiple servers. We also deploy 3 blockchains. Each blockchain contains 4 nodes and uses the Proof-of-Authority consensus algorithm. The block time is set to be 5 seconds. Additionally, the gas calculation method in the experiments is the same as that used in Ethereum. For every experiment, we select Blockchain 1 (bc1) as the execution chain in the IntegrateX system, with the other chains as invoked chains.
(a) Latency
(b) Gas consumption
Figure 5: Latency and gas consumption in the Train-and-Hotel use case.
VII-B Experimental Results
VII-B1 Performance under Train-and-Hotel Use Case
We now compare the latency and gas consumption during the Train-and-Hotel CCSCI process for IntegrateX and GPACT.
As shown in Figure 5(a), IntegrateX decreases total latency by 51.2% compared to GPACT. On bc1 (execution chain), the latency of IntegrateX is reduced by 46.2% compared to GPACT. The reason is that, due to the integrated execution, IntegrateX could integrate and execute all the related execution logic within one block, avoiding multiple rounds of cross-chain execution during CCSCI. On bc2, IntegrateX also reduces the latency by 60% compared to GPACT. This is because the train contract is invoked multiple times, requiring multiple rounds of state transfers in GPACT, whereas in IntegrateX, only a single round of state transfer is needed.
The results in Figure 5(b) show that the total gas consumption of IntegrateX and GPACT is nearly identical in the Train-and-Hotel use case. On bc1, the gas consumption in IntegrateX is larger than GPACT (41.9%). This is mainly because in IntegrateX, the execution chain (bc1) handles all the logic during CCSCI, leading to a higher gas. However, on the invoked chains (bc2 and bc3), IntegrateX achieves lower gas cost compared to GPACT (35.4%, 35.2%, respectively). The main reason is that in IntegrateX, the invoked chains only need to lock and update states without executing any logic.
VII-B2 Performance under Different Invocation Complexity
We evaluate the performance of latency and gas consumption of IntegrateX and GPACT with different CCSCI complexity. We vary the call tree depth to represent different levels of complexity in cross-chain invocations.
(a) Latency
(b) Gas consumption
Figure 6: Latency and gas consumption under different call tree depths.
The latency results are shown in Figure 6(a). Due to the atomic integrated execution mechanism, the latency of IntegrateX and IntegrateX without transaction aggregation (TA) remain stable at around 35 seconds. The latency of GPACT, however, increases linearly as the complexity of CCSCI grows. Specifically, when the call tree depth is 4, IntegrateX significantly reduces latency by 61.2% compared to GPACT. Furthermore, as can be inferred, with invocation complexity further increases, IntegrateX will show greater improvements compared to GPACT.
The gas consumption results are shown in Figure 6(b). Due to the lower invocation complexity, the gas savings from transaction aggregation are not significant, when the call tree depth is 3, the gas consumption for CCSCI in IntegrateX is nearly the same compared to GPACT. As the invocation complexity increases, when the call tree depth reaches 4, transaction aggregation reduces gas consumption in IntegrateX by 9.8% compared to GPACT. And with further increases in invocation complexity, transaction aggregation will reduces more gas consumption in IntegrateX. We also evaluate the gas consumption between IntegrateX and IntegrateX without transaction aggregation. The experimental results show that as the complexity of CCSCI increases, the transaction aggregation mechanism significantly reduces gas consumption in CCSCI. When the call tree depth is 3, the transaction aggregation mechanism reduces gas consumption by 5.4%, and when the call tree depth is 4, it reduces gas consumption by 19.3%. This suggests that the transaction aggregation mechanism achieves greater improvements in gas consumption as the complexity of CCSCI increases.
VII-B3 Concurrency Performance with Fine-Grained State Lock
We now measure the concurrency performance of IntegrateX with and without the fine-grained state lock mechanism. To show the concurrency performance, we initiate multiple cross-chain calls on the same state simultaneously and compare the time consumed by the cross-chain dApp. The experimental results are shown in the Figure 7.
Figure 7: Latency of handling CCSCI in IntegrateX with/without Fine-Grained State Lock (FGSL).
The experimental results show that with fine-grained state lock mechanism, the latency of IntegrateX remains stable at around 35 seconds regardless of the number of simultaneous calls. This demonstrates that the fine-grained state lock can maintain good transaction concurrency. On the other side, as the number of simultaneous calls increases, the latency increases linearly in IntegrateX without fine-grained state lock. When the number of simultaneous invocations reached 6, the average latency of the IntegrateX system rises to 178.8 seconds, representing an increase of approximately 400.1%.
TABLE I: Gas consumption during cross-chain deployment.
Contract Type w/o LSD (Gas) w/ LSD (Gas)
Train Contract 1,459,626 749,383
Hotel Contract 1,524,151 389,383
VII-B4 Logic-State Decoupling
We evaluate the gas consumption during the cross-chain smart contract deployment, using the Train-and-Hotel contracts as an example. We compare the results of the train and hotel contracts with and without logic-state decoupling. The experimental results in Table I indicate that, after applying logic-state decoupling, gas consumption during off-chain clone and deployment is reduced by 48.6% for the train contract and 74.5% for the hotel contract. The main reason is that, the train contract is implemented with more complex logic, while the hotel contract stores more states. As can been seen, for contracts with substantial state but relatively simple logic, logic-state decoupling can significantly reduce gas consumption.
VIII Discussions and Limitations
Potential Application Scenarios. Beyond the example of the train-and-hotel problem mentioned in this paper, IntegrateX can be extended to a wide range of application scenarios. One promising use case is cross-chain flash loans [54]. Flash loans are atomic, uncollateralized lending protocols that allow users to borrow funds at nearly zero cost, perform other operations, and then repay the loan. However, these processes require the guarantee of overall atomicity, meaning that either all steps succeed or they all fail. Due to this requirement for atomicity, existing flash loan protocols are limited to intra-chain operations. With IntegrateX, which provides overall atomicity for cross-chain dApps, users can efficiently perform cross-chain flash loan operations. Other promising application scenarios include, but are not limited to, cross-chain atomic arbitrage and cross-chain supply chain management.
Learning Cost for Developers. The logic-state decoupling and fine-grained state lock mechanisms may slightly increase the development learning curve for smart contract developers. Fortunately, we have proposed a set of guidelines to assist developers, and the mechanisms are flexible (as discussed in Section IV-A and V-C), allowing developers to freely decide whether to implement them. Furthermore, we can provide formal documentation, SDKs, and other resources (following existing standards such as Wormhole [55] and IBC [17]) to guide developers in secure and efficient development and auditing, thereby reducing the learning curve. Additionally, incentive mechanisms (e.g., token rewards) are widely adopted in the industry to encourage developers to build and utilize our system. In future research, we could even explore AI-based semi-automated smart contract tools to further address this challenge.
Moreover, logic-state decoupling offers several additional benefits. For instance, it facilitates modular programming principles in smart contract development, which helps reduce subsequent upgrade and maintenance costs while improving contract security. When an issue arises in a specific contract module, developers can conduct targeted audits and resolve vulnerabilities efficiently.
Support for Heterogeneous Chains. IntegrateX currently supports blockchains that run different consensus protocols but share the same smart contract execution environment. As mentioned in the paper, to ensure cross-chain transaction security across blockchains with different consensus protocols, IntegrateX waits until consensus on the source chain is finalized (or highly likely to be finalized) before committing the cross-chain transaction to the target chain. Additionally, while this paper focuses on IntegrateX’s implementation on EVM-compatible blockchains, it can also be modified to operate between non-EVM-compatible blockchains that share the same smart contract execution environment.
However, a current limitation of IntegrateX is that it cannot operate between blockchains with different smart contract execution environments. Fortunately, this issue could potentially be addressed using advanced techniques such as code virtualization [56]. Expanding IntegrateX to support integrated execution across blockchains with different smart contract environments is a future research direction we aim to explore.
Trade-off Between Flexibility and Load Balancing. In IntegrateX, cross-chain dApp providers have the flexibility to select any chain as the execution chain for integrated execution, based on their preferences. However, this flexibility may introduce a potential issue: if many cross-chain dApp providers choose the same chain as the execution chain, that chain could become a hotspot, potentially degrading its performance. One possible solution is for a third party (e.g., IntegrateX) to manage load balancing by selecting the execution chain on behalf of the cross-chain dApp providers. However, this approach could introduce centralization risks. Additionally, as discussed in the paper, developers’ choice of which chain to run their dApps on often involves considerations beyond performance, such as ecosystem compatibility and business partnerships. How to better achieve load balancing and how to strike a trade-off between performance and flexibility are important questions that warrant future research.
Mitigating Malicious Application Layer Components. In public blockchain scenarios, there are common strategies to mitigate malicious behavior from application layer components (e.g., dApp providers, users). For instance, a malicious cross-chain dApp provider might attempt to maliciously lock certain states to prevent their usage by others. Such behavior can be countered using contract-based authorization or blacklisting mechanisms (widely used in existing dApp development [57]). For example, an intra-chain dApp provider can pre-arrange with a cross-chain dApp provider and authorize trusted cross-chain dApp providers (through their associated addresses) in their contracts, allowing only authorized cross-chain dApp providers to invoke and lock their states. Similarly, intra-chain dApp providers can blacklist specific cross-chain dApp providers in their contracts to block their interactions. Additionally, gas fee mechanisms can serve as a deterrent to malicious application layer components attempting to launch flooding attacks against the blockchain.
Cross-Chain vs. Cross-Shard. Some existing works have explored the issue of cross-shard smart contract handling [58, 51]. However, cross-chain and cross-shard scenarios are fundamentally different, and their solutions cannot be directly applied to cross-chain contexts. The primary reasons are as follows: First, research on blockchain sharding typically involves modifications to the underlying system. In contrast, a key requirement of cross-chain protocols or systems is that they must not require modifications to the underlying blockchains, ensuring better compatibility with existing blockchain systems. Second, a blockchain sharding system usually has a beacon chain responsible for coordinating progress across shards, which is impractical in cross-chain protocols. After all, in cross-chain scenarios, each blockchain essentially belongs to a different system. Finally, in cross-chain contexts, blockchains are often heterogeneous (e.g., different consensus protocols), which is uncommon in blockchain sharding systems.
Inter-Chain Shared Security. In IntegrateX, each blockchain is assumed to have a proportion of malicious nodes lower than its fault tolerance threshold (i.e., they are secure). This is a widely accepted assumption in most existing works. However, recent research has begun exploring how to achieve secure cross-chain interoperability protocols in scenarios where individual blockchains may not be secure, by sharing security across multiple blockchains [47]. IntegrateX could adopt similar ideas through modifications to achieve shared security among blockchains. However, how to design and implement inter-chain shared security within IntegrateX while still maintaining efficiency is an important direction for future research.
IX Conclusion
In this paper, we introduce IntegrateX, a system that enhances cross-chain interoperability by ensuring overall atomicity and efficiency in cross-chain smart contract invocations across EVM-compatible blockchains. The proposed hybrid cross-chain smart contract deployment protocol enables the logic contract to be cloned and deployed onto a single blockchain efficiently and securely. The proposed cross-chain smart contract integrated execution protocol allows efficient intra-chain integrated execution for cross-chain smart contract invocations while guarantee overall atomicity. Experimental results show that, compared to the state-of-the-art baseline, IntegrateX is able to reduce a significant portion of latency while maintaining low gas cost and high concurrency, particularly excelling in complex cross-chain interactions.